OpenStudy (anonymous):

Determine whether (6,-2) iw a solution of the system. x-y=8 2x+3y=6

OpenStudy (dumbcow):

6-(-2) = 6+2 = 8 True 2*6 + 3*-2 = 12 -6 = 6 True